The second term of a geometric series is 3 and the common ratio is 4/ 5 . Find the sum of first 23 consecutive terms of this series. 


1
Expert's answer
2021-06-21T08:01:12-0400

Consider the geometric series "a, ar, ar^2, ..."


"a_n=ar^{n-1}, n=1, 2, ..."

"S_n=\dfrac{a(1-r^n)}{1-r}, r\not=1"

Given "r=\dfrac{4}{5}, ar=3."

Find "a"


"a(\dfrac{4}{5})=3=>a=\dfrac{15}{4}"

"S_{23}=\dfrac{\dfrac{15}{4}(1-(\dfrac{4}{5})^{23})}{1-\dfrac{4}{5}}"

"S_{23}=\dfrac{75}{4}\big(1-(\dfrac{4}{5})^{23}\big)"